# Quality Time: Carbon-Aware Quality Adaptation for Energy-Intensive Services

### v1

@@ Abstract.
The energy demand of modern cloud services, particularly those related to generative AI, is increasing at an unprecedented pace. While hyperscalers are collectively failing to meet their self-imposed emission reduction targets, they face increasing pressure from environmental sustainability reporting across many jurisdictions. To date, carbon-aware computing strategies have primarily focused on batch process scheduling or geo-distributed load balancing. However, such approaches are not applicable to services that require constant availability at specific locations, due to latency, privacy, data, or infrastructure constraints.
In this paper, we explore how the carbon footprint of energy-intensive services can be reduced, by adjusting the fraction of requests served by different service quality tiers. We show, that by adapting the the quality of responses with respect to local carbon intensity, we can achieve additional carbon savings beyond resource and energy efficiency. Building on this, we introduce a multi-horizon optimization, that reaches close-to-optimal carbon savings under realistic conditions, and can dynamically adapt the service quality for best-effort users to stay within an annual carbon budget. Our approach can reduce the emissions of large-scale LLM services, which we estimate at multiple 10,000 tons of CO2 annually, by up to 10 %.
Carbon-aware computing, quality of service, LLM inference, generative AI, service-oriented computing
submissionid: ¡PAPER ID¿ccs: Social and professional topics Sustainabilityccs: Networks Cloud computingPreprint v1. Under review.
@@ 1. Introduction
Operating modern cloud services can require substantial computing resources and energy, which is contributing directly to the rapidly growing carbon footprint of computing systems (S&P Globl Market Intelligence, 2023). Alongside traditional resource-intensive services like real-time data analytics, video processing, recommendation systems, and streaming (particularly for cloud gaming and virtual reality), the growing use and high energy demand of generative AI is currently driving the expansion of data centers at a high pace (Bashir et al., 2024; Wu et al., 2024). As a result, major cloud providers, who collectively pledged to reduce their carbon footprint to zero by 2030, have been increasing their emissions significantly in recent years (Google, 2024; Microsoft, 2024; Meta Platforms, 2023) and are expected to row back on sustainability commitments (Lawrence, 2024).
@@
Besides their social responsibility to minimize environmental impact, large companies are increasingly under pressure to report their greenhouse gas (GHG) emissions as part of environmental sustainability reporting (Project, 2024; initiative, SBTi). While cloud providers generally report Scope 1 emissions (direct emissions from company-owned or controlled sources, such as on-site fuel combustion), reporting on Scope 2 emissions (i.e. electricity consumed by data centers) is becoming mandatory in many parts of the world (Deloitte, 2024; European Parliament and Council of the European Union, 2022; UK Department for Business, Energy & Industrial Strategy, 2019; Japan Ministry of Enconomy, Trade and Industry (2023), METI). Moreover, upcoming policy changes will mandate that large companies also disclose Scope 3 emissions (European Parliament and Council of the European Union, 2022; Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023), which include all indirect emissions that occur across the value chain. These include emissions from the use of of cloud services, which will affect not only service providers but also customers of energy-intensive services.
@@
The concept of carbon-aware computing has emerged as a response to these challenges: Carbon-aware strategies aim to reduce emissions beyond energy efficiency (i.e., reducing demand) by additionally taking into account how carbon-intensive the consumption of electricity at a current time and location is. However, existing carbon-aware approaches primarily focus on determining the ideal time and location of executing flexible batch processing tasks (Wiesner et al., 2021; Radovanovic et al., 2022; Lechowicz et al., 2023; Zheng et al., 2020; Wiesner et al., 2024; Hanafy et al., 2024; Sukprasert et al., 2024b). Services, on the other hand, usually require permanent availability and have little to no delay-tolerance, which is why carbon efficiency in services has been mainly explored in the context of geo-spatial load shifting (Gsteiger et al., 2024; Murillo et al., 2024; Souza et al., 2023; Zhou et al., 2016, 2013). Yet, many services cannot simply be distributed across regions due to latency constraints (i.e., proximity of servers and users), data availability, privacy regulations, security concerns, and—most importantly—infrastructure limitations. For instance, generative AI inference typically relies on high-performance GPUs or other accelerators, that are available only at specific locations.
@@
Figure 2. We investigate how adjusting the QoR with respect to carbon intensity can reduce emissions beyond savings related to mere energy efficiency.
@@
In this work, we focus on the temporal variability of carbon intensity (see Figure 1) and explain how adapting service quality with respect to this metric does not only reduce energy demand, but can yield additional carbon savings. We define the quality of responses (QoR) as a type of quality of service (QoS) metric, that describes the proportion of requests that are served by different service quality tiers. For instance, machine learning models can be deployed in smaller, quantized, or pruned versions that use fewer resources and consume less energy, but compromise on response quality (Liang et al., 2021). Similarly, in LLM inference, adjusting the number of returned tokens is another way to directly impact the energy demand of requests (Li et al., 2024b).
Adjusting service quality is common practice for avoiding overload (Meza et al., 2023; Park et al., 2024), but has also recently been proposed as a strategy for improving service sustainability (Vitali, 2022). Figure 2 illustrates, how carbon-aware QoR adaptions can reduce emissions by an additional 5-10 % in many regions—beyond the savings related to reduced energy demand.
Contributions
@@
Towards more carbon-efficient cloud services, we make the following contributions in this paper:
@@
    We model emissions for data center operators and cloud tenants (Scope 2 and 3 reporting) and show that both perspectives are equivalent while optimizing for energy and carbon efficiency.
@@
    We propose a forecast-based multi-horizon optimization for minimizing service emissions, which can dynamically adjust the QoR for best-effort users to stay within a predefined annual carbon budget.
@@
    We evaluate the upper-bound potential and practicability of carbon-aware QoR adaptation by simulating a large-scale LLM inference service across two request traces in ten different regions.
All results from this analysis are fully reproducible. Code and data will be made publicly available upon acceptance.
@@ 2. Background and Motivation
We review the current state of GHG reporting for data centers and cloud services, and motivate our work by outlining anticipated regulatory changes.
2.1. Energy-Intensive Cloud Services
The growing computational demand of generative AI is a major factor for the expansion of data centers worldwide (S&P Globl Market Intelligence, 2023). In particular, the rapid adoption of large language models (LLMs) requires immense processing power, not only for initial training but also for inference at scale (Faiz et al., 2024; Luccioni et al., 2024). As generative AI evolves towards multi-modality, emerging formats like video are expected to further increase electricity demand in the future (Li et al., 2024a).
Beyond generative AI, a range of current and emerging services are contributing to increasing energy demands. Video streaming, predominantly in high-resolution formats, now accounts for 82 % of all internet traffic (Systems, 2024). Platforms such as cloud gaming, virtual reality, and the metaverse require real-time graphics rendering, which consumes significant power (Stoll et al., 2022; Mills et al., 2019; Kattakinda et al., 2024). Additionally, scientific simulations—including climate modeling (Acosta et al., 2024), drug discovery (Wynendaele et al., 2021), and computational fields like astrophysics (Taffoni et al., 2019), fluid dynamics (Yang et al., 2024), and bioinformatics (Grealey et al., 2022)—are growing in complexity and increasingly integrate AI inference services.
2.2. Carbon Footprint of Cloud Providers
Despite previous commitments to achieve fully carbon-free or even ‘carbon-negative’ operations by 2030, Google has reported a 48 % increase in GHG emissions in the last five years (Google, 2024), Microsoft has announced a 29.1 % rise in emissions since 2020 (Microsoft, 2024), while Meta increased their emissions by 46 % in 2022 alone (Meta Platforms, 2023)—predominantly because of the training and deployment of generative AI. The rapid expansion of data centers, combined with the slow adoption of clean energy and efficiency measures, suggests that reaching net-zero emissions in the near future is highly unlikely (Lawrence, 2024).
@@ 2.2. Carbon Footprint of Cloud Providers
2.3. Growing Pressure from GHG Reporting
The GHG Protocol (Protocol, 2001) is a globally recognized framework that provides standards for organizations to report their GHG emissions. It forms the basis for many national and international climate programs, such as the Carbon Disclosure Project (CDP) (Project, 2024) and Science Based Targets initiative (SBTi) (initiative, SBTi) and categorizes emissions into three scopes:
  • Scope 1: Direct GHG emissions that occur from sources that are owned or controlled by the company.
  • Scope 2: Indirect GHG emissions from the generation of purchased electricity consumed by the company.
  • Scope 3: All other indirect emissions that occur across the value chain, such as from suppliers or services.
For the IT sector, Scope 2 and 3 emissions are the most relevant categories. Their reporting is becoming increasingly standardized and mandatory across major jurisdictions (Deloitte, 2024; European Parliament and Council of the European Union, 2022; UK Department for Business, Energy & Industrial Strategy, 2019; Japan Ministry of Enconomy, Trade and Industry (2023), METI; Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023), with significant implications for data center operators, cloud tenants, and—ultimately—users of energy-intensive services.
In the US, the Securities and Exchange Commission (SEC) is advancing regulations that require public companies to fully disclose their Scope 2 emissions (Deloitte, 2024). While, the disclosure of Scope 3 emissions stays voluntarily, for now, California soon requires large companies to also report Scope 3 emissions (Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023). Meanwhile, the EU’s Corporate Sustainability Reporting Directive (CSRD) mandates Scope 2 and 3 disclosures for large companies (European Parliament and Council of the European Union, 2022) beginning in 2024, aligning with the EU’s broader climate goals under the European Green Deal (European Commission, 2019). From 2026 on, this will also include non-EU companies with significant operations in the EU.
Although current legislation around the world mostly addresses Scope 1 and 2 emissions, the significance of Scope 3 emissions has been emphasized repeatedly (Hettler and Graf‐Vlachy, 2024; Hertwich and Wood, 2018; Radonjič and Tompa, 2018). Therefore, also jurisdictions outside the EU and California are likely to require Scope 3 emissions reporting in the future.
2.4. Implications for Energy-Intensive Cloud Services
Figure 3. Service providers either operate their own infrastructure or rent from cloud providers. Either way, electricity usage will be reflected in GHG reporting, as operational carbon emissions are passed on to cloud tenants. Data center Scope 3 emissions are often reffered to as embodied emissions.
When reporting emissions from the perspective of providers of energy-intensive services, we must distinguish between two cases, as illustrated in Figure 3:
@@ 2.4. Implications for Energy-Intensive Cloud Services
    They operate data centers themselves, in which case the operational carbon footprint is mainly determined by electricity purchases, i.e. Scope 2 emissions.
@@
    They rent infrastructure from cloud providers, in which case the data center’s Scope 2 and 3 emissions will be passed down the value chain, which is an active topic of research (Amaral et al., 2023; Choochotkaew et al., 2023; Hè et al., 2023; Schmidt et al., 2023; Westerhof et al., 2023; Han et al., 2024).
Today’s, landscape is mixed: While, for example, Google relies on its proprietary infrastructure to train and serve AI models like Gemini (Pichai and Hassabis, 2024), Microsoft Azure provides the computational resources for operating OpenAI’s ChatGPT (OpenAI, 2023).
While we will address the two cases in our problem formalization, both are affected by stricter GHG reporting, as outlined earlier. Following, the increasing emissions of computing and the growing pressure for comprehensive accounting will soon have a direct impact on service cost. Carbon pricing mechanisms covered 23 % of global emissions in 2023, where revenues reached a record \$104 billion, and there is a clear trend towards stricter rules and higher prices (World Bank, 2024).
2.5. Assumptions on Future Regulatory Policies
As the effectiveness of yearly and globally valid energy certificates is highly questionable (Bjørn et al., 2022; Holzapfel et al., 2023; Bashir et al., 2023), it is likely that future regulatory policies will demand time-based and location-based reporting of operational emissions. Based on current initiatives from industry (Energy, 2023; Cramer and Goodman, 2023) and legislation (Becker, 2021), we assume that operational carbon emissions of electricity usage will be reported hourly—or even sub-hourly (Haugen et al., 2024; Bashir et al., 2023).
We also follow the predominant assumption in related works that future regulations will use average carbon intensity (ACI), reported in carbon-equivalent GHG per unit of energy (gCO2/kWh), as a fundamental metric to determine the operational carbon emissions at specific times and locations (Gsteiger et al., 2024; Wiesner et al., 2021; Hanafy et al., 2024; Sukprasert et al., 2024b; Radovanovic et al., 2022). ACI is defined as a weighted average of all produced energy in a region, weighted by the carbon intensity of the energy source. Some organizations, like Microsoft (Buchanan et al., 2023) advocate the use of marginal carbon intensity (MCI), which refers to the additional carbon emissions generated from meeting incremental demand at a given time. Although the two metrics can result in different scheduling decisions (Sukprasert et al., 2024a), both can be used in our framework. In our analysis, however, we assume ACI, as MCI is very ambiguous to compute in practice (Wiesner et al., 2021) and therefore a lot less likely to make it into official policies.
@@ 3. Carbon-Aware QoR Adaptation
Offering diverse service quality tiers for different user groups is common practice in the operation of cloud services. For example, content delivery networks provide faster and more reliable service to higher-tier users, while AI platforms like OpenAI and Gemini reserve access to larger and more advanced models for paying customers. Furthermore, service providers already degrade service quality during periods of high load to prevent system overload (Park et al., 2024; Meza et al., 2023). Such measures often target best-effort users: For instance, free-tier users of LLM inference services can experience highly increased latencies at certain times, to not endanger the QoS of subscribed users.
Recently, there have been proposals to design applications that can intentionally sacrifice QoS, not only to manage economical interests or load peaks, but as a strategy to improve sustainability (Vitali, 2022; Vitali et al., 2023). Especially in AI inference, we have numerous options for fine-grained quality tiering that have different impacts on service quality and energy efficiency. For instance, service quality tiers for LLMs can range from entirely different models (e.g. number of parameters), quantization (many models existin in variants from 4bit integers to 32bit floating point numbers), as well as number of returned tokens (i.e. we generate shorter replies for some users), which can have a significant impact on the carbon footprint of inference (Li et al., 2024b).
In this work, we investigate the effect of adapting service QoR over time for improved carbon efficiency within a single region, i.e., without geo-distributed load balancing.

### v2

@@ Abstract.
The energy demand of modern cloud services, particularly those related to generative AI, is increasing at an unprecedented pace. While hyperscalers collectively fail to meet their self-imposed emission reduction targets, they face increasing pressure from environmental sustainability reporting across many jurisdictions. To date, carbon-aware computing strategies have primarily focused on batch process scheduling or geo-distributed load balancing. However, such approaches are not applicable to services that require constant availability at specific locations due to latency, privacy, data, or infrastructure constraints.
In this paper, we explore how the carbon footprint of energy-intensive services can be reduced by adjusting the fraction of requests served by different service quality tiers. We show that adapting this quality of responses with respect to grid carbon intensity can lead to additional carbon savings beyond resource and energy efficiency. Building on this, we introduce a forecast-based multi-horizon optimization that reaches close-to-optimal carbon savings and is able to automatically adapt service quality for best-effort users to stay within an annual carbon budget. Our approach can reduce the emissions of large-scale LLM services, which we estimate at multiple 10,000 tons of CO2 annually, by up to 10 %.
Sustainable computing, quality of service, LLM inference, green AI
submissionid: ¡PAPER ID¿ccs: Social and professional topics Sustainabilityccs: Networks Cloud computingPreprint v2. Under review.
@@ 1. Introduction
Operating modern cloud services can require substantial computing resources and energy, which directly contributes to the rapidly growing carbon footprint of computing systems (S&P Globl Market Intelligence, 2023). Alongside traditional resource-intensive services like real-time data analytics, video processing, recommendation systems, and streaming, the growing use and high energy demand of generative AI is currently driving the expansion of data centers at a high pace (Bashir et al., 2024; Wu et al., 2024). As a result, major cloud providers, who collectively pledged to reduce their carbon footprint to zero by 2030, have been increasing their emissions significantly in recent years (Google, 2024; Microsoft, 2024; Meta Platforms, 2023) and are expected to row back on sustainability commitments (Lawrence, 2024).
@@
Besides their social responsibility to minimize environmental impact, large companies are increasingly under pressure to report their greenhouse gas (GHG) emissions as part of environmental sustainability reporting (Project, 2024; initiative, SBTi). While reporting on Scope 2 emissions (GHG emissions caused by electricity consumption) is becoming mandatory in many parts of the world (Deloitte, 2024; European Parliament and Council of the European Union, 2022; UK Department for Business, Energy & Industrial Strategy, 2019; Japan Ministry of Enconomy, Trade and Industry (2023), METI), upcoming policy changes will mandate that large companies also disclose Scope 3 emissions (European Parliament and Council of the European Union, 2022; Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023)—including emissions from the use of cloud services. This will affect not only service providers, but also users of energy-intensive services.
@@
The concept of carbon-aware computing has emerged as a response to such challenges: Carbon-aware strategies aim to reduce emissions beyond energy efficiency by additionally taking into account how carbon-intensive the consumption of electricity at a current time and location is. Especially in the context of scheduling flexible batch workloads, the potential of temporal and spatial load shifting is well understood (Wiesner et al., 2021; Radovanovic et al., 2022; Lechowicz et al., 2023; Zheng et al., 2020; Wiesner et al., 2024; Hanafy et al., 2024b; Sukprasert et al., 2024b; Hanafy et al., 2024a). However, for continuously running services, where individual requests have little to no delay-tolerance, existing approaches exclusively focus on geographical load balancing (Gsteiger et al., 2024; Murillo et al., 2024; Souza et al., 2023; Zhou et al., 2016, 2013). However, many services cannot simply be distributed across regions due to data locality requirements, regulatory constraints, privacy and security concerns, or infrastructure limitations. For instance, generative AI inference typically relies on high-performance GPUs or specialized accelerators, which are often only available at specific locations. Carbon-aware approaches for services that are constrained to a single region remain largely unexplored.
@@
Figure 2. We investigate how adjusting the QoR with respect to carbon intensity can reduce emissions beyond savings related to reduced energy demand.
@@
In this work, we focus on the temporal variability of carbon intensity (see Figure 1) and explain how adapting service quality with respect to this metric does not only impact energy demand, but can yield additional carbon savings. We define the quality of responses (QoR) as a type of quality of service (QoS) metric that describes the proportion of requests served by different service quality tiers. For instance, machine learning models can be deployed in smaller, quantized, or pruned versions that use fewer resources and consume less energy, but compromise on response quality (Liang et al., 2021). Our approach minimizes the operational carbon emissions of energy-intensive services by optimizing resource provisioning and load balancing across service quality tiers, while ensuring compliance with service level objectives (SLOs) for different user groups.
Figure 2 shows the accumulated energy usage and carbon emissions for an example scenario with two quality tiers over one week. Serving all requests at the high quality tier (QoR = 1) uses roughly twice the energy compared to serving them at the low quality tier (QoR = 0). Balancing requests evenly between the two tiers results in intermediate energy usage. We explore how strategically timing quality reductions based on carbon intensity can further reduce emissions by performing a comprehensive analysis across different regions, request patterns, and QoR constraints. Based on this, we propose an approach that achieves close-to-optimal savings despite real-life challenges such as forecast inaccuracies and the high computational complexity of the underlying optimization.
@@
Contributions.  Towards more carbon-efficient cloud services, we make the following contributions:
@@
    We model emissions from the perspective of data center operators and cloud tenants (Scope 2 and 3 reporting) and prove that both models lead to equivalent decisions.
@@
    We propose a forecast-based multi-horizon optimization for minimizing service emissions, which can automatically adjust the QoR for best-effort users to stay within a predefined annual carbon budget.
@@
    We evaluate the upper-bound potential and practicability of carbon-aware QoR adaptation by simulating a large-scale LLM inference service on eight real and synthetic request traces in ten different regions.
All experiments in this paper are fully reproducible. All code and data will be made publicly available upon acceptance.
@@ 2. Background and Motivation
Despite previous commitments to achieve fully carbon-free or even “carbon-negative” operations by 2030, Microsoft recently announced a 29.1 % rise in emissions since 2020 (Microsoft, 2024), Google reported a 48 % increase in GHG emissions within five years (Google, 2024), while Meta increased their emissions by 46 % in 2022 alone (Meta Platforms, 2023)—predominantly because of the training and deployment of generative AI. The rapid expansion of data centers, combined with the slow adoption of clean energy and efficiency measures, suggests that reaching net-zero emissions in the near future is highly unlikely (Lawrence, 2024).
@@
Energy-intensive cloud services.  The rapid adoption of large language models (LLMs) requires immense processing power, not only for training but also for inference at scale (Faiz et al., 2024; Luccioni et al., 2024). As generative AI evolves towards multi-modality, emerging formats like video are expected to further increase future electricity demand (Li et al., 2024a).
Beyond generative AI, a range of current and emerging services are contributing to increasing energy demands. Video streaming, predominantly in high-resolution formats, now accounts for 82 % of all internet traffic (Systems, 2024). Platforms such as cloud gaming, virtual reality, and the metaverse require real-time graphics rendering, which consumes significant power (Stoll et al., 2022; Mills et al., 2019; Kattakinda et al., 2024). Additionally, scientific simulations such as climate modeling (Acosta et al., 2024), drug discovery (Wynendaele et al., 2021), astrophysics (Taffoni et al., 2019), fluid dynamics (Yang et al., 2024), and bioinformatics (Grealey et al., 2022) are growing in complexity and increasingly integrate AI services.
Growing pressure from GHG reporting.  The reporting of Scope 2 emissions, i.e. indirect GHG emissions caused by consuming electricity, is becoming increasingly standardized and mandatory across major jurisdictions (Deloitte, 2024; European Parliament and Council of the European Union, 2022; UK Department for Business, Energy & Industrial Strategy, 2019; Japan Ministry of Enconomy, Trade and Industry (2023), METI; Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023), with significant implications for data center operators, cloud tenants, and—ultimately—users of energy-intensive services.
In the US, the Securities and Exchange Commission (SEC) is advancing regulations that require public companies to fully disclose their Scope 2 emissions (Deloitte, 2024). While, the disclosure of Scope 3 emissions—i.e. all indirect emissions that occur across the value chain—stays voluntarily, for now, California soon requires large companies to also report Scope 3 emissions (Paul, Weiss, Rifkind, Wharton & Garrison LLP, 2023). Meanwhile, the EU’s Corporate Sustainability Reporting Directive (CSRD) mandates Scope 2 and 3 disclosures for large companies (European Parliament and Council of the European Union, 2022) beginning in 2024, aligning with the EU’s broader climate goals under the European Green Deal (European Commission, 2019). From 2026 on, this will also include non-EU companies with significant operations in the EU. As the significance of Scope 3 emissions has been emphasized repeatedly (Hettler and Graf‐Vlachy, 2024; Hertwich and Wood, 2018; Radonjič and Tompa, 2018), also jurisdictions outside the EU and California are likely to require Scope 3 emissions reporting in the future.
Implications for energy-intensive services.  When reporting emissions from the perspective of cloud service providers, we must distinguish between two cases:
@@
    They operate data centers themselves, in which case the operational carbon footprint is mainly determined by electricity usage, i.e. Scope 2 emissions.
@@
    They rent infrastructure from cloud providers, in which case the data center’s emissions will be passed down the value chain als Scope 3 emissions, which is an active topic of research (Amaral et al., 2023; Choochotkaew et al., 2023; Hè et al., 2023; Schmidt et al., 2023; Westerhof et al., 2023; Han et al., 2024).
Today’s, landscape is mixed: While, for example, Google relies on its proprietary infrastructure to train and serve AI models like Gemini (Pichai and Hassabis, 2024), Microsoft Azure provides the computational resources for operating OpenAI’s ChatGPT (OpenAI, 2023). We will address the two cases in our problem formalization, however, both are affected by stricter GHG reporting. Following, the increasing emissions of computing and the growing pressure for comprehensive accounting will soon have a direct impact on service cost. Carbon pricing mechanisms covered 23 % of global emissions in 2023, where revenues reached a record \$104 billion, and there is a clear trend towards stricter rules and higher prices (World Bank, 2024).
Assumptions on future regulatory policies.  As the effectiveness of annual and globally valid energy certificates is highly questionable (Bjørn et al., 2022; Holzapfel et al., 2023; Bashir et al., 2023), it is likely that future regulatory policies will demand time-based and location-based reporting of operational emissions. Based on current initiatives from industry (Energy, 2023; Cramer and Goodman, 2023) and legislation (Becker, 2021), we assume that operational carbon emissions of electricity usage will be reported hourly—or even sub-hourly (Haugen et al., 2024; Bashir et al., 2023).
We also follow the predominant assumption in related works that future regulations will use average carbon intensity (ACI), reported in carbon-equivalent GHG per unit of energy (gCO2/kWh), as a fundamental metric to determine the operational carbon emissions at specific times and locations (Gsteiger et al., 2024; Wiesner et al., 2021; Hanafy et al., 2024b; Sukprasert et al., 2024b; Radovanovic et al., 2022). ACI is defined as a weighted average of all produced energy in a region, weighted by the carbon intensity of the energy source. Some organizations, like Microsoft (Buchanan et al., 2023) advocate the use of marginal carbon intensity (MCI), which refers to the additional carbon emissions generated from meeting incremental demand at a given time. Although the two metrics can result in different scheduling decisions (Sukprasert et al., 2024a), both can be used in our framework. In our analysis, however, we assume ACI, as MCI is very ambiguous to compute in practice (Wiesner et al., 2021) and therefore a lot less likely to make it into official policies.
@@ 3. Carbon-Aware QoR Adaptation
Offering different service quality tiers is common practice in the operation of cloud services. For example, content delivery networks provide faster and more reliable service to higher-tier users, while AI platforms like OpenAI and Gemini reserve access to larger and more advanced models for paying customers. Furthermore, service providers already degrade service quality during periods of high load to prevent system overload (Park et al., 2024; Meza et al., 2023). Such measures often target best-effort users: For instance, free-tier users of LLM inference services can experience highly increased latencies at certain times, to not endanger the QoS of subscribed users.
Recently, there have been proposals to design applications that can intentionally sacrifice QoS, not only to manage economical interests or load peaks, but as a strategy to improve sustainability (Vitali, 2022; Vitali et al., 2023). Especially in AI inference, we have numerous options for fine-grained quality tiering that have different impacts on service quality and energy efficiency. For instance, service quality tiers for LLMs can vary based on factors such as the choice of model (e.g., number of parameters), quantization (ranging from 4-bit integers to 32-bit floating-point), and the number of tokens generated (i.e. providing shorter replies (Li et al., 2024b)), and can have a significant impact on the carbon footprint of inference.
In this work, we examine the impact of dynamically adjusting service QoR over time to enhance carbon efficiency within a single region, without relying on geo-distributed load balancing.